A service delivered to users through a cloud-hosted web server has a target response time of 200 ... A service delivered to users through a cloud-hosted web server has a target response time of 200 milliseconds (0.2 of a second). The latency between London and Dublin is 50 milliseconds and the latency between London and Frankfurt is 60 milliseconds. What percentage increase in the amount of time for processing the response is achieved by a business in London if it switches from a web-server in Frankfurt to one in Dublin?

give the number rounded to the nearest whole number %
Answer #1

Latency between London and Dublin connection = 50 ms

Latency between London and Frankfurt connection = 60 ms

Switching the server from Frankfurt to one in Dublin will lead to reduce the latency by (60-50) = 10 ms

Percentage increase in amount of time for processing the response = 10 ms/60 ms * 100 = 1/6 * 100 = 16.67%

Hence the closest whole number answer will be equal to 17%

Note - Post any doubts/queries in comments section.